OpenStudy (anonymous):

simplify the equation by combining like terms 6x-3x=9

OpenStudy (anonymous):

i keep getting 3

OpenStudy (anonymous):

Combine the x terms to get 3x = 9 divide by 3 to get x=3

OpenStudy (anonymous):

6x-3x=9 3x=9 3x/3=9/3 x=3